OpenStudy (anonymous):

Find the limit. lim t^2-36 / t-6 x->6

OpenStudy (anonymous):

i put "does not exist" as my answer but that's wrong

OpenStudy (anonymous):

it becomes 0/0 so shouldn't that be DNE

OpenStudy (anonymous):

factor t^2-36 / t-6 t^2-36 / t-6 =(t-6)(t+6)/(t-6) =(t+6) then evaluate lim (t+6) x->6

OpenStudy (anonymous):

simplify it becomes lmt t+ 6 as t > 6 lmt = 12

OpenStudy (anonymous):

ooooh i see

OpenStudy (anonymous):

good